4 .A random sample of the birth weights of 24 babies has a mean of 3103 grams and a standard deviation of 696 grams. Construct a 95% confidence interval estimate of the mean birth weight for all such babies.

if possible include graph
Answer #1

Solution :

Given that,

\bar x = 3103

s = 696

n =24

Degrees of freedom = df = n - 1 =24 - 1 = 23

a ) At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

  \alpha/ 2= 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,90 = 2.069 ( using student t table)

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.069* ( 696 / \sqrt 24)

= 293.9437

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

3103 - 293.9437< \mu < 3103+ 293.9437

2809.0563 < \mu < 3396.9437

( 2809.0563 ,3396.9437)